OpenStudy (deadshot):

What is the exact value for \[\sin (- \frac{ \pi }{ 3 })\] ?

OpenStudy (luigi0210):

Here's a rule: \[\LARGE sin(-x)=-sin(x)\]

OpenStudy (deadshot):

so then it would become \[-\sin (\frac{ \pi }{ 3 })\]

OpenStudy (luigi0210):

Yea :)

OpenStudy (deadshot):

and since \[\frac{ \pi }{ 3 }\] = 60 degrees, it would become \[-\sin (60)\] right?

OpenStudy (luigi0210):

Yup

OpenStudy (deadshot):

so would the answer be \[- \frac{ \sqrt{3} }{ 2 }\]?

OpenStudy (luigi0210):

*confetti* There ya go

OpenStudy (deadshot):

Thanks! :)

OpenStudy (luigi0210):

You're welcome